economics question be clear, show steps and correct solution for rating solve in 30 minutes for thumbs up rating

A piece of land may be bought for $600,000 to be strip-mined for coal. Annual net income will be $500,000 a year for 10 years. At the end of the 10 years, the surface of the land will be restored as required by a federal law on strip mining. The reclamation will cost $1,500,000 more than the resale value of the land after it is restored. Using a 10% interest rate compounded quarterly, find the present worth of this project (Don't use the $ sign in your answer)
Answer #1

Rate per quarter =10%/4 =2.5%
EAR=(1+2.5%)^4-1=10.381289%
Number of Periods =10

Present Worth of the project =PV of Annual Net Income -PV of Reclamation cost -Initial Investment
=500000*((1-(1+10.381289%)^-10)/10.381289%)-1500000/(1+10.381289%)^10-600000=1,863,952.57